A firm has established a distribution network for the supply of a raw material critical to its manufacturing. Currently there are two origins for this material, which must be shipped to three manufacturing plants. The current network has the following characteristics:

Costs

Plant 1

Plant 2

Plant 3

Supply

Raw material source 1

$6

$8

$9

400

Raw material source 2

$4

$7

$3

600

Demand

500

500

500

Supply < Demand

The firm has identified two potential sites for a third raw material source; these are identified as Candidate A and Candidate B. From A, the costs to ship would be $9 to Plant 1, $10 to Plant 2, and $12 to Plant 3. From B, these costs would be $11, $14, and $8. The new source, wherever it is located, will have a capacity of 500 units. Solve with the transportation method. Which site should be selected?
Answer #1

Cost with candidate A = Sum of (Cost of unit * Number of units) = =6*400+4*100+10*500+3*500 = $9300

Solution below:

Costs Raw material source 1 Raw material source 2 Candidate A Plant 1 Plant 2Plant 3 Supply 400 600 500 Supply = Demand $4 $1

Cost with candidate B= Sum of (Cost of unit * Number of units) = =6*400+4*100+14*500+3*500 = $11300

However, as we note in the solution we are supplying 500 units from Source 2 and Candidate B to Plant 3 and Plant 2 respectively. The cost per unit = 14+3 = $17. If we interchange the source to the plant the cost is 7+8 = 15

Hence, Optimal cost with Candidate B ==6*400+4*100+8*500+7*500 = $10300

Solution below:

Costs Raw material source 1 Raw material source 2 Candidate B uppl 54 $11 57 $14 400 600 500 Supply Demand Demand ldentify th

Candidate A should be selected at a lower cost.